"The ratio of my age to my mother's age today is 3:5.
In five years time the ratio will be 7:11.
Let x be the sum of our ages today.
Set up a linear equation and use it to find our ages in five years.
Start the equation like this:
3x/8+5 =..."
I have started the linear equation but do not know how to continue it. This is all I have so far:
3x/8+5 = 7/18(x+10)
Thank You

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

"The ratio of my age to my mother's age today is 3:5.
In five years time the ratio will be 7:11.
---
Your age now: 3x
Mother's age now: 5x
---
Equation:
(3x+5)/(5x+5) = 7/11
---
11(3x+5) = 7(5x+5)
33x + 55 = 35x + 35
2x = 20
x = 10
---
Now Ages:
Your age now = 3x = 30
Mother's age now = 5x = 50
==================================
Cheers,
Stan H.

Answer by solver91311(24713)



I don't see where the sum of the ages is very helpful. You need two variables, one for each person's age.

Let represent the child's age and let represent the mothers age.

Situation Today:



Which is to say:



Then you can say:



Situation five years hence:



Which is to say:



Cross-multiply:



Substitute for and then solve for . Then calculate to derive .

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism